UT Jun 8 2022 - DEC Night



Night Info

  • Sunset 19:26
  • Nautical Twilight Ends 20:29
  • Astro Twilight Ends 21:07
  • Night Center 00:19
  • Moon sets 01:24
  • Astro Twilight Starts 03:31
  • Nautical Twilight Starts 04:08
  • Sunrise 05:11

Staff

  • OSA: D.G.Huerta
  • ISp: C. Bolyard
  • Mtn Manager: R.Hansen
  • ISA/Observer: B.Rothberg/J.Power
  • AO: J.Power

Night Plan

  • Starting configuration: LBTI
  • DONE LBTI - PhaseCam engineering and Eng tasks - twilight-00:14hr
  • Reconfigure for LUCIs 00:14-00:20
  • ESM characterization 00:20-01:20
    • DONE FF1456+3337 - mixed mode ESM and SL- visible twilight-03hr
      • Bin 4 ESM (R=15.7)
      • Scripts: /home/lbto/Service_Observing/2022A_D_E/MODS/Rothberg/midzUlirgs/midz_ULIRGs_V19.xml
        • seeing up to 1.2", thin clouds
        • 1hr
  • DONE LUCI-1 Full AO Tests 01:25-02:25
    • set z0 tests w/Bin 1 continued
    • OT: /home/lbto/Service_Observing/2022A_D_E/LUCI/AO_ESM/AO_Testing_V90.xml
  • ESM Characterization 02:30-03:30
    • Dual BLR - mixed mode ESM and SL - visible all night
      • Bin 2 ESM (R=12.3)
      • Scripts: /home/lbto/Service_Observing/2022A_D_E/LUCI/DualBLR/SDSSJ1713+3256_ESM_SL.xml
        • seeing up to 1.2", thin clouds
        • 1hr

Poor Weather Backup

  • Bino-LUCIs
    • Local Mergers - Rothberg
      • Conditions: If you can guide you can observe:
      • UGC 9829 - RA 15hr Dec -01 - visible twilight-02hr
      • 1.5hr-2hr depending on conditions
      • OT: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V50.xml
    • ICRF3 project
      • OT: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/ICRF3/ICRF3_LUCI_V25_LBTO.xml
      • Pick based on visibility and priority
    • Photometric Standards
      • Must be clear - any seeing
      • If seeing so terrible can't guide, then export as track mode
      • OT: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards/LUCI_Phot_ZP_OT_V4.xml
    • ESM testing in poor conditions
      • Can be run to characterize performance in poor conditions (contined testing)
      • Runs through K,H,J - dithers in and out of FOV
      • Parent Directory: /home/lbto/Service_Observing/2022A_D_E/LUCI/AO_ESM/ESM
        • Mixed mode scripts have already been generated and updated to do mixed mode
        • M92 - bright (R=9.4) and faint (R=15.2)- 1hr each - visible all night
        • M52 - R=10.7 = visible 01hr-twilight
  • Bino-MODS
    • Local Mergers - Rothberg
      • Conditions - Up to 2" seeing, 1-2 mags extinction
      • IRAS 21219-1757 - 2.5hr w/overheads
      • OT: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V50.xml
      • Scripts made for any combo of missing MODS channels
      • visible 02:30hr-twlight - observe SpecStd before - get as many photons as can before twilight too bright
    • Spectrophotometric Standard in all configuations
      • Must be clear, any seeing

Night Log

Closed Dome

  • UT 01:59 - bringing up LUCI GUIs - init all, test darks
    • L1 0002-0004 - 2.51sec dark
    • L2 0003-0004 - 2.51 sec dark
    • all OK

On-Sky

LBTI
  • UT 02:39 - enclosure open - LBTI starting up
  • UT 04:22 - LBTI handing over to LBTO early
Bino-LUCIs
ESM Imaging - FF1106+3201
  • UT 04:24 - reconfiguring for bino-LUCIs
  • UT 04:32 - putting N30fieldstop away in LUCI-1 while we are at 90el and 341rot
  • UT 04:34 - moving to pointing star
    • ran into copointing error
  • UT 04:40 - moving to pointing star - clouds have thinned out more on all-sky
    • swing and a miss
    • DX struggling with pointing
  • UT 04:50 - still working on DX side pointing (LBTI was showing 40" offsets at start of night)
  • UT 04:54 - sending collination preset - 0.67" GCSL, 0.78" GCSR
  • UT 04:58 - sending preset
    • 27.3 degrees from Moon Bin 4 - test of Bin 4 near the Moon too!
    • accidentally sent bino preset, slip of the click, all fixed
  • UT 05:02 - preset resent
    • AO Ref Rmag=15.6, WFSmag = 17.3 (surprisingly fainter than expected, thought it would be 16.8) - showing 3.6photons/subapp, on offsets 0.9 photons/subapp
  • UT 05:08 - closed loop, J.Power increasing tip-tilt (pasued after preset in script)
  • UT 05:09 starting exposures
    • streak through the frame in one of the DIT
    • AOREF 0.47" L1, 0.60" L2
    • Fieldstop on LUCI-2 is a bit off - not worth the time to adjust as target is not close to edge
    • Frame #0010 - 0.45" L1, 0.54" L2
    • Frame #0011 - 0.36" L1, 0.46" L2, ellipticities are lower for L1
    • Frame #0013 - 0.40" L1, 0.43" L2, ellipticities are 0.03 vs 0.24 - Guider is 0.94" on DX
    • Frame #0015 - 0.38" L1, 0.60" L2 - ellipticities are 0.08 vs 0.22
    • Frame #0019 - 0.45" L1, 0.57" L2 -
    • Frame #0021 - 0.40" L1, 0.43" L2 - ellipticities 0.07 vs 0.21
    • El 35deeg, DX guider shows 0.59" seeing - also good test of ESM Bin 4 faint end, in Moonlight in low elevations...
    • Frame #0031 - 0.41" L1, 0.52" L2 - ellipticiities 0.03 vs 0.06 (el 32...really into the muck)
    • UT 06:00 - DX guider shows 0.75" at 32El
    • L1 0005-0034
    • L2 0005-0034
LUCI-1 AO Testing (Bin 1)
HD 151670 Bin 1 (R=9.0) - Optimize Gain, no GOPT
  • UT 06:01 - moving to higher elevation
  • UT 06:03 - manually putting N30 fieldstop into place at 341 rot and EL 90 while re-authorizing for LUCI-1 only
  • UT 06:08 - sending preset - subframe to 128pix needed
  • UT 06:12 - pupils are grossly misaligned
  • UT 06:13 - J.Power interrupted before we had a RIP
  • UT 06:14 - pupils needed to be manually bumped to align them
  • UT 06:15 - camera lens railed in X - we might be ok to continue? - we are centered so shouldn't cause further issues - LUCI-1 instrument rotator = 283
  • UT 06:17 - taking time for real pupils and pupil mask difference to get closer (should be 0.1 or less)
  • UT 06:18 - going to run optimize but this will certainly cause issues
  • UT 06:20 - currently ramping gains - pupils show 2 dark and 2 light diagonal from each other
  • L1 0035 - see a double image - taking a 2nd exposure to see if this is really a double - and I see double again This likely explains the issues - Angular separation is 0.29"
  • HD151760_Double.png
  • Rmag=9.0, WFSmag = 10.
BD+37 2879 Bin 1 (R=8.3) - Optimize Gain, No GOPT, No NCPA
  • UT 06:26 - preset sent
  • Rmag=8.3, WFSmag=9.9
  • 128 pix sub-frame centered on 1133,1157
  • L1 0037-acq
  • L1 0038 - confirm OK
  • my dithers of +/1" are too big for the 1.92" wide - generated new script with +/-0.5" (two cubes might be unusable as a result)
  • UT 06:41 - seems more stable without GOpt - running through full script
  • L1 0039-0053
BD+37 2879 Bin 1 (R=8.3) - Optimize Gain, GOpt on, No NCPA
  • UT 06:50 - will reopitmize and apply optical gain
  • UT 06:53 - starting exposures
  • watching on GEIRS looks like it might be some elongation in PSF
  • L1 0054-0068
BD+37 2879 Bin 1 (R=8.3) - Optimize Gain, GOpt on, NCPA On
  • UT 07:05 - optimizing and turning on NCPA
  • UT 07:08 - starting exposures
  • UT 07:09 - relaunching GEIRS1 window - it got stuck displaying - useful to watch cubes live - 1-2min downtime
  • Had to resend subframe to LUCI - but when I did that (after pausing script) that seemed to trigger it to continues in script - i don't recall clicking GO
  • Gopt looking less noisy,
  • UT 07:14 - 10% safeskip - PSF started falling apart
  • UT 07:17 - frame-to-frame the PSF looks worse now
  • L1 0079 - 24% safeskips - can see this in the PSF
  • L1 0080 - 50% safeskips
  • L1 0081 - 50% safeskips
  • UT 07:19 - turning off NCPA and will repeat three exposures without dithering
  • UT 07:22 - taking TNs
  • UT 07:22 - reoptimizing, which resets optical gain
  • UT 07:24 - At EL 84 now - transit is at 85deg elevation
  • UT 07:25 - loops opened - re-closing loop
  • UT 07:28 - safeskipping again while re-optimizing
  • UT 07:29 - re-optimizing - first solution seemed not acceptable
  • UT 07:31 - reset everything to test NCPA, already safe-skipping and open loop - haven't taken any LUCI exposures
  • UT 07:32 - 92% safeskip - AdSec Arbitrator offline
  • UT 07:34 - need to resend the preset after addressing the offline issue
  • UT 07:34 - resetting GEIRS to full frame
  • L1 0069-0081
BD+37 2879 Bin 1 (R=8.3) - Optimize Gain, GOpt on, No NCPA
  • UT 07:35 - resending preset
  • SX guider before loop close 0.77"-0.84"
  • UT 07:40 - safeskipping
  • UT 07:42 - loop closed
  • L1 0081-Acq
  • L1 0082-0087 - NCPA off
  • UT 07:47 - turning NCPA on, not re-optimzing
  • UT 07:49 -taking exposures
  • NCPA is driving optical gain down
  • UT 07:51 - loop opened
  • UT 07:51 - re-optimizing
  • UT 07:55 - applying NCPA again
  • UT 07:57 - cirrus rolled back in again
  • UT 07:58 - loops opened AGAIN! - curled edge appeared on the mirror GUI at 10:30
  • made it four exposures (16sec each) before the loop opoened
  • L1 0088-0095 NCPA On
  • Conclusion: With optical gain - worked well, throw in NCPA and hilarity ensures
Bino-LUCIs
ESM Imaging - FF1456
  • UT 08:01 - authorizing for bino luci
  • UT 08:01 - moving LUCI-1 rotaotor to 341 to put N30fielstop mask away, El 90
  • UT 08:06 - sending preset
  • UT 08:80 - no star found, restoring IE CA
  • UT 08:10 - doing pointing check
  • UT 08:14 - resending preset
  • wispy cirrus rolled through, can see more waves of it approaching from the west
  • UT 08:19 - saying 0 for tip-tilt - J.Power working on that
  • UT 08:20 - loops closed, taking exposures
  • UT 08:39 - seeing 1.1
  • UT 08:39 - IQ looks like L2 > L1 0 J.Power bumping down gains
  • L1 0108- 0.55", L2 0047 - 0.64"
  • UT 08:46 - DX guider shows 0.77"
  • UT 08:50 - changed H01 but that seemed to make it worse, resetting back
  • UT 08:52 - OCAM not live - offset did not complete - loops are paused
    • Looks like we have to resend the preset
  • Rmag =15.7, WFSMag = 17.2
    • L1 - 0096-0106 - gains might be too high
    • L1 - 0107-0113
    • L2 - 0035-0052
  • UT 08:58 - moved L2 fieldstop a bit right (20pix)
  • UT 08:59 -J. Power resending preset - but also hit reset - lets see what happens
  • UT 09:01 - DX guider shows 0.75"-0.8"
  • Resuming from where it was paused and OCAM went not-live
  • UT 09:17 - down to 2.4 photons/subapp WFSMag=17.4 - 17.8
  • thin cirrus rolling in - might explain why L2 is again showing better IQ - last 30min see 0.8 mag drop in relative flux on guiders
  • Again, found the limits of ESM...
  • UT 09:31 - clouds are thickening, although seeing is <1" the flux loss is getting more and more - will switch to poor weather science with MODS when this script finishes
    • L1 0114-0132
    • L2 0053-0071
    • This was a good EXTREME ESM comparison w/SL - and also can eek something out thats useful
  • UT 09:37 - script done
  • UT 09:38 - putting LUCIs in safe config, J.Power stopping AO
Bino- MODS - Poor Weather Backup
IR21219-1757 - LS Dual Grating
  • UT 09:39 - reconfiguring for MODS
  • UT 09:53 Pointing and collimating
  • UT 10:08 Sending Acquisition Preset to IRAS21219-1757
    • Average seeing 0.83 on teh left and 0.74 ont eh right guiders
    • MODS2:

      Computed Slit Alignment Offset:
      dX = 3.436 arcsec
      dY = 8.647 arcsec

      MODS2 Offset Command:
      offsetxy 3.436 8.647 rel

    • mods1:

      Target selection done.

      Computed Slit Alignment Offset:
      dX = -0.600 arcsec
      dY = 9.665 arcsec

      MODS1 Offset Command:
      offsetxy -0.600 9.665 rel. (sent y -9 by accident so y +2 to recover that mistake)

  • UT10:23 sending science
    • band of clouds in the south west (where we are pointed) about half a mag average extinction. Seeing avaera 0.75" on the right and 0.77 on the left on the guiders
    • mods[1/2]r.20220608.0006-7 (second exposures are short)
      mods[1/2]b.20220608.0003-4 (second exposures are short)
    • 10:35 Seeing deteriorating a bit on the guiders now and DIMM, 1" on the guider and 1.2" on the DIMM
    • 10:58. Seeing 1.35" on the DIMM, 1.05 average on the right and 1.11" average on teh left guider. thin cloud
    • Stopped and readout at just over 5 min so we can get the standard on the second exposure
    • mods1r readout lag
  • 11:05 Sending preset for Feige110
    • MODS1

      Computed Slit Alignment Offset:
      dX = 0.007 arcsec
      dY = 12.186 arcsec

      MODS1 Offset Command:
      offsetxy 0.007 12.186 rel
      additional

      Computed Slit Alignment Offset:
      dX = -0.381 arcsec
      dY = -0.241 arcsec

      MODS1 Offset Command:
      offsetxy -0.381 -0.241 rel

    • mods2

      Computed Slit Alignment Offset:
      dX = 4.692 arcsec
      dY = 8.672 arcsec

      MODS2 Offset Command:
      offsetxy 4.692 8.672 rel
      additional:

      Computed Slit Alignment Offset:
      dX = -0.627 arcsec
      dY = -0.294 arcsec

      MODS2 Offset Command:
      offsetxy -0.627 -0.294 rel

  • 11:11 Sending science script for Feige 110
    • mods[1/2]r.20220608.0011-13
      mods[1/2]b.20220608.0005-7
    • Seeing 1.19" average on teh right and 1.12" average on the left.
    • mods1r readout lag again
11:25 Closing

11:41 Starting HBand Flats
  • luci1.20220608.0133-142, luci2.20220608.0072-81 (probe may not have been parked, david parking)
  • luci1.20220608.0143-152, luci2.20220608.0082-91

-- BarryRothberg - 03 Jun 2022
Topic revision: r7 - 08 Jun 2022, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback